OpenStudy (anonymous):

write a polynomial function in standard form with the given zeros. x=3,0

OpenStudy (anonymous):

y = (x-a)(x-b) is a polynomal with zeros x=a,b

OpenStudy (anonymous):

helps ?

OpenStudy (jdoe0001):

the idea been x = 3 => x-3 = 0 => (x-3) = 0 x =0 => x - 0 = => (x-0) = 0 thus now, but using the "zeros" or "solutions" of the quadratic equation, that is the "x-intercepts" we know its roots, that is (x -3) and (x-0) so just multiply those to get your original polynomial

OpenStudy (anonymous):

multiply wat

OpenStudy (jdoe0001):

\(\bf x = 3 \implies x-3 = 0 \implies \color{red}{(x-3)} = 0\\ \quad \\ x =0 \implies x - 0 =0 \implies \color{red}{(x-0)} = 0\\ \quad \\ \color{red}{(x-3)(x-0)} = 0\implies \square x^2+\square x+\square = 0\)

OpenStudy (anonymous):

i dont know how to get the answer tho

OpenStudy (jdoe0001):

just multiply the binomials, the exercise assume you do know polynomials multiplication

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=ZMLFfTX615w

OpenStudy (anonymous):

thank u

OpenStudy (jdoe0001):

yw